What are two numbers whose quotient is -6 and whose sum is -12?
Tried -12 and 2---- 12/2=-6 but (-12) + 2= -10
Tried -18 and 3------ no to both parts

Answer by ptaylor(2198)

Let x=one number
and Let y=the other number
Now we are told that:
x/y=-6----eq1
and x+y=-12----eq2
In eq1, multiply each side by y and we get:
x=-6y---eq1a----substitute this into eq2
-6y+y=-12
-5y=-12
y=12/5--the other number
x=-6y=-6(12/5)=-72/5=the first number
CK
(-72/5)/(12/5)=-6; (-72/5)(5/12)=-6 ok
(12/5)+(-72/5)=-12
-60/5=-12
-12=-12----ok
Hope this helps---ptaylor